A rectangular garden is 16 feet wide. Its length is 7 feet more than twice the width. What is the area of the garden?
Answer by hamsanash1981@gmail.com(151)

Given breath = 16 , then length = 2(16) +7 = 39
Area of a rectangular garden = 39*16 = 624 square feet.